OpenStudy (anonymous):

A rectangle has sides of length 2x+4 and 6. A) find another set of dimensions for a rectangle that has the same perimeter. b) Find another set of dimensions for a rectangle that has the same area.

OpenStudy (dumbcow):

A) P = (2x+4) + (2x+4) +6 +6 = 4x+20 as long as the dimensions add up to 4x+20 they work for example a square with sides of (x+5) would have same perimeter B) A = (2x+4)*6 = 12x+24 as long as length*width = 12x+24 they work for example i can factor out a 2 12x+24 = 2(6x+12) so length =2 width = 6x+12